On a sunny day, a building and its shadow form the sides of a right triangle. If the hypotenuse is 34 m long and the shadow is 24 m, how tall is the building?
Answer by mananth(16946)

the leg of the right triangle 1s 24 & hypotenuse is 34 meters
using the pythagorus theorem
leg1^2 +leg2^2 = hyp ^2
leg1^2 + 24^2 = 34^2
leg1^2 = 34^2 - 24^2
leg1^2 = 1156-576
leg1^2=580
leg1 = sqrt 580
leg1 = 24.08 which is the height of the building
Ananth